OpenStudy (anonymous):

Find the height of a square pyramid that has a volume of 75 cubic feet and a base length of 5 feet. 5 feet 9 feet 25 feet 45 feet

OpenStudy (anonymous):

the formula for a square pyramid is v=1/3 b^2h you already have the volume and the base so you just plug it in 75=1/3(5)^2h 75=1/3(25)h 75=8.325h 9.oo9009009=h your answer is 9

OpenStudy (anonymous):

9feet 75/25=3 3*(3/1)=9 h=9